OpenStudy (anonymous):

solve the formula for the indicated variable. 5x-4y=7 for y can you help me to understand?

OpenStudy (anonymous):

y =( 5/4)x-7/4

OpenStudy (anonymous):

just keep one varaible and transfer the other

OpenStudy (diyadiya):

4y = 12 - 5x y = ( 12 - 5x) / 4 y = (12 / 4) - (5x / 4) y = 3 - (5x / 4)

OpenStudy (anonymous):

thanks